Solve the system by the addition method:
6x+3y=51
2x-6y=38
Thank you

Answer by robertb(5830)

Multiply the top equation by 2: 12x + 6y = 102.
Add corresponding sides with the bottom equation: 2x-6y = 38;
14x = 140, or x = 10.
Then multiply the bottom eqn by -3: -6x+18y = -114.
Addc corresponding sides with the top equation: 6x + 3y = 51.
21y = -63, or y = -3.